Compsre the bond lengths of : N-N, O-O,   C-C and F-F single bonds. Please give a proper explanation for each bond length! 

Grade:11

3 Answers

Arun
5 years ago

In chemistry there is nothing like always.You have to see any element characteristics than decide its property. the case you are asking about

1)Noteworthy exceptions are single bonds between the period 2 atoms of groups 15, 16, and 17 (i.e., N, O, F), which are unusually weak compared with single bonds between their larger congeners. It is likely that the N–N, O–O, and F–F single bonds are weaker than might be expected due to strong repulsive interactions between lone pairs of electrons on adjacent atoms

2)N-N>F-F ,in F there are greater number of lone pairs than in N so it experiences strong repulsion between its electrons that make it vulnerable to be broken

3)F-F>O-O it's so because F atom is very small that enables strong electrostatic attraction between nucleus and electrons

 

Bond Lengths

C-C = 1.54 A

O- O = 1.48A

N-N = 1.47 A

Tony
5 years ago
Also compare C-C, N-N AND O-O Simultaneously plzz! I will surely then approve  your answer! Can you tell me some trick to compare them easily using just 1 or 2 properties! 
Mahima Kanawat
5 years ago
Dear student 
Bond length - single bond>double bond >triple bond
Therefore C-C, F-F > O-O > N-N
Now comparing C-C and F-F, fluorine is smaller in size than C.. Therefore, C-C > F-F
WITH REGARDS
MAHIMA 
ASKIITIANS FORUM EXPERT